# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 7. Rank and Genus of countably generated projective modules
defined by $\Psi([P])=[P_{Q}]=[S^{r}]$, where $S$ is a simple $\Lambda_{Q}$-module. Then we have a monoid morphism, that we also call $\Psi$
@@
defined by $\Psi([P])=r$ if $P$ is finitely generated and $P_{Q}\cong S^{r}$ and $\Psi([P])=\infty$ if $P$ is not finitely generated. We call $\Psi([P])$ the rank of $P$.
@@
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})}$
defined by $\Phi([P])=([P_{\mathfrak{m}}])_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}$. Recall that two countably generated projective right $\Lambda$-modules $P$ and $P^{\prime}$ have the same genus if and only if $\Phi([P])=\Phi([P^{\prime}])$. We shall refer to $\Phi$ as the genus morphism.
In this section, we want to describe the image of $\Psi$ and of $\Phi$.
At this point, we have plenty of information about $V^{*}(\Lambda)$. By Theorem 6.16, $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, by Lemma 6.6 any ideal of $\mathcal{T}(\Lambda)$ is a trace ideal of a finitely generated projective module and, by Corollary 6.10, $\mathcal{T}(\Lambda)$ is finite.
In general, there may be plenty of finitely generated projective right modules over $\Lambda$ with the same genus that are not isomorphic. For example, let $R=\Lambda$ be a Dedekind domain. Then any localization of a projective module at a maximal ideal is free. So two projective modules have the same genus if and only if they have the same rank. Certainly, two finitely generated projective modules over a Dedekind domain with the same rank are not, in general, isomorphic.
The situation changes for projective $\Lambda$-modules whose isomorphism class is in $B(\Lambda)$.
@@
Let $\Lambda$ be a locally semiperfect, torsion-free algebra over an $h$-local domain $R$ such that $\Lambda_{Q}$ is simple artinian. Then $\Phi(B(\Lambda))\subseteq\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}B(\Lambda_{\mathfrak{m}})$. Moreover, if $[P]$ and $[P^{\prime}]$ are two elements in $B(\Lambda)$, then $\Phi[P]=\Phi[P^{\prime}]$ if and only if $[P]=[P^{\prime}]$.
Proof.  For this proof, we are going to use the notation of Corollary 2.7 that is, an element $[P]\in B(\Lambda)$ can be thought of as the isomorphism class of a pair $(P/PI,I)$ where $I\in\mathcal{T}(\Lambda)$ and $P/PI$ is a finitely generated projective $\Lambda/I$-module. By Proposition 6.11, for each maximal ideal $\mathfrak{m}$ of $R$, $[P_{\mathfrak{m}}]\in B(\Lambda_{\mathfrak{m}})$ and $I_{\mathfrak{m}}$ is the ideal minimal with respect to the property that $P_{\mathfrak{m}}/P_{\mathfrak{m}}J$ is finitely generated.
Hence, $\Phi([(P/PI,I)])=([(P_{\mathfrak{m}}/P_{\mathfrak{m}}I_{\mathfrak{m}},I_{\mathfrak{m}})])_{\mathfrak{m}\in\mathrm{maxSpec}\,(R)}.$ By Lemma 1.14, it follows that $\Phi$ is injective restricted to $B(\Lambda)$.
The following straightened version of the Package Deal Theorem to our setting will be the key ingredient to describe $\Phi(V^{*}(\Lambda))=\Phi(V(\Lambda)\sqcup B(\Lambda)$.
@@
Proof.  Let $k$ be such that $\Lambda_{Q}\cong S^{k}$. By Lemma 6.8, there is a finite set of maximal ideals of $R$, $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$, such that $\Lambda_{\mathfrak{m}}\cong M_{k}(T_{\mathfrak{m}})$ for some local ring $T_{\mathfrak{m}}$ whenever $\mathfrak{m}$ is a maximal ideal not in $A=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$. Note that if $\mathfrak{m}\not\in A$ then $\Lambda_{\mathfrak{m}}$ is isomorphic to a ring of $k\times k$ matrices over a local ring, so it has a unique indecomposable projective module $X(\mathfrak{m})$ such that $X(\mathfrak{m})_{Q}\cong S$ and any projective $\Lambda(\mathfrak{m})$-module is isomorphic to a suitable number of copies of $X(\mathfrak{m})$. Therefore, for any $\mathfrak{m}\not\in A$, $P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
Let $q$ be a multiple of $k$ bigger than $r$ and let $F=\Lambda^{q/k}$. Further let $U$ be a finitely generated submodule of $F$ such that $U_{Q}\cong S^{r}$. Since the embedding $U\hookrightarrow F$ splits over $\Lambda_{Q}$, it splits over almost all maximal ideals of $R$ [AHP, Lemma 4.2]. Then $U$ is a finitely generated $\Lambda$-module which is torsion-free as an $R$-module, such that $U_{\mathfrak{m}}$ is projective for almost all maximal ideals $\mathfrak{m}$ of $R$ and $U_{Q}\cong S^{r}$. So let $B$ be a finite set of maximal ideals of $R$ such that $A\subseteq B$, and satisfying that $U_{\mathfrak{m}}$ is projective for any $\mathfrak{m}\not\in B$. In particular, for any $\mathfrak{m}\not\in B$, $U_{\mathfrak{m}}\cong P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
By Lemma 5.6, for any $\mathfrak{m}\in B$ we can choose $U(\mathfrak{m})$ a submodule of $U_{\mathfrak{m}}$ such that $U(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{m}\not\in B$ we set $U(\mathfrak{m})=U_{\mathfrak{m}}$.
Note that for any maximal ideal $\mathfrak{m}$ of $R$, $U(\mathfrak{m})_{Q}=U_{Q}$ since all these modules have length $r$ over $\Lambda_{Q}$. Therefore, we can apply Theorem 5.9 and conclude that there exists a finitely generated $\Lambda$-module $P$ of $U$ such that $P(\mathfrak{m})\cong U(\mathfrak{m})=P_{\mathfrak{m}}$ for every $\mathfrak{m}\in\operatorname{mSpec}R$. By Corollary 5.4, $P$ is a projective module over $\Lambda$.
Now we introduce the notation needed to explain in detail how the image of the genus map $\Phi$ and of the rank map $\Psi$ is.
Notation 7.3.
@@
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})\cong\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}{\mathbb{N}}_{0}^{*}}$
@@
For $i=1,\dots,\ell$ let $U_{i,1},\dots,U_{i,t_{i}}$ be representatives of indecomposable projective right $\Lambda_{\mathfrak{m}_{i}}$-modules, and let $r_{i,1},\dots,r_{i,t_{i}}$ be the lengths of their localizations over $\Lambda_{Q}$, i.e., $(U_{i,j})_{Q}\cong S^{r_{i,j}}$.
Note that for any $i=1,\dots,\ell$, if $\Lambda_{\mathfrak{m}_{i}}$ is isomorphic, as a right $\Lambda_{\mathfrak{m}_{i}}$-module, to $U_{i,1}^{m_{i,1}}\oplus\dots\oplus U_{i,t_{i}}^{m_{i,t_{i}}}$, then we should have the relation
@@
Now we construct examples of locally semiperfect algebras over an $h$-local domain with an infinite number of maximal ideals, showing there are no other restrictions on the parameters $k,\ell,t_{i},r_{ij},m_{ij}$ than the $\ell$ equations above.
@@
    $\Lambda$ is a subring of ${\rm M}_{k}(\mathbb{Q})$ and $\Lambda_{Q}={\rm M}_{k}(\mathbb{Q})$.
@@
From $(iii)$ it follows that $\Lambda/J$ is a direct sum of simple modules. Hence $\Lambda/J$ is semisimple artinian and therefore $J=J(\Lambda)$. Also note that the canonical projection $e_{i,j,j}\Lambda\to e_{i,j,j}\Lambda/e_{i,j,j}J$ is a projective cover of $e_{i,j,j}\Lambda/e_{i,j,j}J$. So $\Lambda$ is semiperfect and $e_{i,j,j}\Lambda$ are indecomposable projectives. Also $e_{i,a,a}\Lambda\cong e_{j,b,b}\Lambda$ if and only if $i=j$. So we have that $\Lambda=\bigoplus_{i=1}^{t}\bigoplus_{a=1}^{m_{i}}e_{i,a,a}\Lambda\cong\bigoplus_{i=1}^{t}(e_{i,1,1}\Lambda)^{m_{i}}$. Also it is easy to check that $e_{i,a,a}{\rm M}_{k}(\mathbb{Q})$ has length $r_{i}$ since $|A_{i,a}|=r_{i}$.
@@
Now the package deal theorem finds a $\mathbb{Z}$-submodule $\Lambda\subseteq{\rm M}_{k}(\mathbb{Z})$ such that $\Lambda_{(p_{i})}=\Lambda(p_{i})$ for $i=1,\dots,\ell$ and $\Lambda_{(q)}={\rm M}_{k}(\mathbb{Z}_{(q)})$ if $q$ is a prime different from $p_{1},\dots,p_{\ell}$. It is easy to see that $\Lambda$ has to be a subring of ${\rm M}_{k}(\mathbb{Z})$ and that $\Lambda_{(p_{i})}\simeq\Lambda(p_{i})$ as rings.
@@
We keep the Notation 7.3. First we observe that
${\Phi\colon V^{*}(\Lambda)}$${\prod_{\mathfrak{m}\in\operatorname{mSpec}(\Lambda)}V^{*}(\Lambda_{\mathfrak{m}})=\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})\times\prod_{\mathfrak{m}\not\in\mathcal{S}}V^{*}(\Lambda_{\mathfrak{m}})}$
So for any $\mathfrak{m}\not\in\mathcal{S}$, $V^{*}(\Lambda_{\mathfrak{m}})\cong{\mathbb{N}}_{0}^{*}$ and, for any $[P]\in V^{*}(\Lambda)$, this isomorphism sends $[P_{\mathfrak{m}}]$ to its rank which is $\Psi[P]$. So the genus map is completely determined by its image on $V^{*}(\Lambda_{\mathfrak{m}})=\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})$.
Then there is a monoid morphism
@@
Where the last isomorphism assigns to $\Phi^{\prime}([P])$ the element
@@
Conversely, by Proposition 7.2, an element $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ corresponds to a finitely generated projective right $\Lambda$-module if and only if there exists $r$ such that the components of $\mathbf{X}$ satisfy the system of equations (10). So this describes all the possible genus of finitely generated projective $\Lambda$-modules.
So we have proved the following.
@@
Let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$ in $\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$. Then $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ is an element of $\mathcal{A}$ if and only if there exists $r\in{\mathbb{N}}_{0}$ such that the components of $\mathbf{X}$ satisfy
@@
If $\ell=1$, and $U_{1},\dots,U_{t}$ are the indecomposable projective modules of $\Lambda_{\mathfrak{m}}$, up to isomorphism, and $\Psi(U_{i})=r_{i}$ for $i=1,\dots,t$. Then $\Psi(V(\Lambda))$ is the submonoid of ${\mathbb{N}}_{0}$ generated by $r_{1},\dots,r_{\ell}$. By Proposition 7.2, $\Phi^{\prime}(V(R))\cong{\mathbb{N}}_{0}^{t}$, because, for any $i=1,\dots,\ell$, there exists an (indecomposable) finitely generated projective right $\Lambda$-module $P_{i}$ such that $(P_{i})_{\mathfrak{m}_{i}}\cong U_{i}$.
@@
$\sum_{j=1}^{t_{i}}r_{i,j}y_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}y_{i,1}\\ \vdots\\ y_{i,t_{i}}\end{pmatrix}=\infty$ (11)
@@
Let $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$. For $i=1,\dots,\ell$ let $S_{i}(\mathbf{Y})=\{j\mid y_{i,j}=\infty\}$. Observe that $\mathbf{Y}$ is a solution of the equations (11) if and only if $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$.
Let $P$ be a countably generated right $\Lambda$-module such that it is not finitely generated. Then $\Psi(P)=\infty$, which implies that, for any $i=1,\dots,\ell$, there exists $j$ such that the indecomposable module $U_{i,j}$ appears infinitely many times in the decomposition of $P_{\mathfrak{m}}$ as a direct sum of indecomposable modules. Therefore, if $\mathbf{Y}=\Phi^{\prime}([P])$ then $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$. So $\mathbf{Y}$ is a solution of the system (11).
@@
For $i=1,\dots,\ell$, let $P^{\prime}(\mathfrak{m}_{i})=\bigoplus_{j\in S_{i}(\mathbf{Y})}U_{i,j}$ and let $r_{i}$ be the rank of $P^{\prime}(\mathfrak{m}_{i})$. Let $r$ be the least common multiple of $r_{1},\dots,r_{\ell},k$. For $i=1,\dots,\ell$, set $P(\mathfrak{m}_{i})=P^{\prime}(\mathfrak{m}_{i})^{\frac{m}{r_{i}}}$ and set $P(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{\frac{r}{k}}$. By Package deal Theorem 5.12, there exists a finitely generated projective $\Lambda$-module $P^{\prime}$ such that $P^{\prime}_{\mathfrak{m}}\cong P(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$.

### v2

@@ 7. Rank and Genus of countably generated projective modules
defined by $\Psi([P])=[P_{Q}]=[S^{r}]$, where $S$ is a simple right $\Lambda_{Q}$-module. Then we have a monoid morphism that, by abuse of notation, we also call $\Psi$
@@
defined by $\Psi([P])=r$ if $P$ is finitely generated and $P_{Q}\cong S^{r}$, and $\Psi([P])=\infty$ if $P$ is not finitely generated. We call $\Psi([P])$ the rank of $P$, and we refer to $\Psi$ as the rank map.
@@
$\Phi\colon V^{*}(\Lambda)$$\prod_{\mathfrak{m}\in\operatorname{mSpec}(R)}V^{*}(\Lambda_{\mathfrak{m}})$}
defined by $\Phi([P])=([P_{\mathfrak{m}}])_{\mathfrak{m}\in\operatorname{mSpec}(R)}$. We shall refer to $\Phi$ as the genus morphism, and to $\Phi([P])$ as the genus of $[P]$.
We say that two countably generated projective right $\Lambda$-modules $P$ and $P^{\prime}$ are in the same genus if $P_{\mathfrak{m}}\cong P^{\prime}_{\mathfrak{m}}$ for every maximal ideal $\mathfrak{m}$ of $R$, that is, if and only if $\Phi([P])=\Phi([P^{\prime}])$.
In this section, we want to describe the images of $\Psi$ and $\Phi$.
At this point, we have plenty of information about $V^{*}(\Lambda)$. By Theorem 6.17, $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, and by Lemma 6.7, any ideal of $\mathcal{T}(\Lambda)$ is a trace ideal of a finitely generated projective module, and by Corollary 6.11, $\mathcal{T}(\Lambda)$ is finite.
In general, there may be plenty of finitely generated projective right modules over $\Lambda$ in the same genus that are not isomorphic. For example, let $R=\Lambda$ be a Dedekind domain. Then any localization of a projective module at a maximal ideal is free. So two projective modules are in the same genus if and only if they have the same rank. Certainly, two finitely generated projective modules over a Dedekind domain with the same rank are not, in general, isomorphic.
The situation changes for projective $\Lambda$-modules whose isomorphism class is in $B(\Lambda)$, or for finitely generated projective $\Lambda$-modules when $\Lambda$ is a suitable algebra over a semilocal domain $R$.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Then $\Phi(B(\Lambda))\subseteq\prod_{\mathfrak{m}\in\operatorname{mSpec}(R)}B(\Lambda_{\mathfrak{m}})$. Moreover, if $[P]$ and $P^{\prime}]$ are two elements in $B(\Lambda)$, then $\Phi([P])=\Phi([P^{\prime}])$ if and only if $P]=[P^{\prime}]$.
Proof. The first part of the proof follows from Corollary 5.4 and Proposition 6.12.
For the second part of the proof, we are going to use the notation of Corollary 2.5 that is, an element $P]\in B(\Lambda)$ can be thought of as the isomorphism class of a pair $(P/PI,I)$ where $I\in\mathcal{T}(\Lambda)\setminus\{0\}$, and $P$ is relatively $I$-big. Hence, $\Phi([(P/PI,I)])=([(P_{\mathfrak{m}}/P_{\mathfrak{m}}I_{\mathfrak{m}},I_{\mathfrak{m}})])_{\mathfrak{m}\in\operatorname{mSpec}(R)}.$
By Lemma 5.9, there exist $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$, such that $\Lambda/I\cong\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$. Hence, $P/PI\cong\prod_{i=1}^{\ell}P_{\mathfrak{m}_{i}}/P_{\mathfrak{m}_{i}}I_{\mathfrak{m}_{i}}$, and $P_{\mathfrak{n}}=P_{\mathfrak{n}}I_{\mathfrak{n}}$ is free and $I_{\mathfrak{n}}=\Lambda_{\mathfrak{n}}$, for any $\mathfrak{n}\in\operatorname{mSpec}(R)\setminus\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$.
By Theorem 5.11, $I$ is uniquely determined by $I_{\mathfrak{m}}$, $\mathfrak{m}\in\operatorname{mSpec}(R)$. Therefore, by Lemma 1.14, it follows that $\Phi$ is injective when restricted to $B(\Lambda)$.
Proposition 7.2.
Let $\Lambda$ be a locally semiperfect, torsion-free algebra over a semilocal domain $R$. Then:
  • (i)
    $\Lambda$ is semilocal;
  • (ii)
    Assume $\Lambda$ is module-finite over $R$, and let $P_{1}$ and $P_{2}$ be two finitely generated projective right $\Lambda$-modules. Then $P_{1}\cong P_{2}$ if and only if $\Phi([P_{1}])=\Phi([P_{2}])$.
Proof. $(i)$ Let $\operatorname{mSpec}(R)=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$. Consider the embedding
$\alpha\colon\Lambda\to\prod_{i\in\{1,\dots,\ell\}}\Lambda_{\mathfrak{m}_{i}}$
induced by taking the localization map at each component. If $\lambda\in\Lambda$ is such that $\alpha(\lambda)$ is invertible, then for each $i\in\{1,\dots,\ell\}$, there exists $s_{i}\in R\setminus\mathfrak{m}_{i}$ and $\mu_{i}\in\Lambda$ such that $\lambda\mu_{i}=s_{i}$. Then $I=\sum_{i=1}^{\ell}s_{i}R$ is not contained in any maximal ideal of $R$, so $I=R$. Since $R=I\subseteq\lambda\Lambda$, we deduce that $\lambda$ is a right invertible element of $\Lambda$. By symmetry, $\lambda$ is an invertible element of $\Lambda$. This shows that $\alpha$ is a local ring homomorphism (the image of an element is invertible if and only if it is an invertible element of $\Lambda$). Since $\prod_{i\in\{1,\dots,\ell\}}\Lambda_{\mathfrak{m}_{i}}$ is a finite product of semiperfect rings, we deduce that $\Lambda$ is semilocal (cf. [libro, Theorem 4.2]).
$(ii)$ This statement is proved in [AHP, Corollary 4.5].
The following straightened version of the Package Deal Theorem to our setting will be the key ingredient to describe the image of the genus map $\Phi$.
@@
Proof. Let $k$ be such that $\Lambda_{Q}\cong S^{k}$. By Lemma 6.9, there is a finite set of maximal ideals of $R$, $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$, such that $\Lambda_{\mathfrak{m}}\cong M_{k}(T_{\mathfrak{m}})$ for some local ring $T_{\mathfrak{m}}$ whenever $\mathfrak{m}$ is a maximal ideal not in $\mathcal{S}=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}\}$. Note that if $\mathfrak{m}\not\in\mathcal{S}$, then $\Lambda_{\mathfrak{m}}$ is isomorphic to a ring of $k\times k$ matrices over a local ring, so it has a unique indecomposable projective module $X(\mathfrak{m})$ such that $X(\mathfrak{m})_{Q}\cong S$ and any projective $\Lambda(\mathfrak{m})$-module is isomorphic to a suitable number of copies of $X(\mathfrak{m})$. Therefore, for any $\mathfrak{m}\not\in\mathcal{S}$, $P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
Let $q$ be a multiple of $k$ that is greater than $r$, and let $F=\Lambda^{q/k}$. Further, let $U$ be a finitely generated submodule of $F$ such that $U_{Q}\cong S^{r}$. Since the embedding $U\hookrightarrow F$ splits over $\Lambda_{Q}$, it splits over almost all maximal ideals of $R$ [AHP, Lemma 4.2]. Then $U$ is a finitely generated $\Lambda$-module which is torsion-free as an $R$-module, such that $U_{\mathfrak{m}}$ is projective for almost all maximal ideals $\mathfrak{m}$ of $R$ and $U_{Q}\cong S^{r}$. So let $B$ be a finite set of maximal ideals of $R$ such that $\mathcal{S}\subseteq B$, and satisfying that $U_{\mathfrak{m}}$ is projective for any $\mathfrak{m}\not\in B$. In particular, for any $\mathfrak{m}\not\in B$, $U_{\mathfrak{m}}\cong P(\mathfrak{m})\cong X(\mathfrak{m})^{r}$.
By Lemma 5.7, for any $\mathfrak{m}\in B$, we can choose $U(\mathfrak{m})$ a submodule of $U_{\mathfrak{m}}$ such that $U(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{m}\not\in B$, we set $U(\mathfrak{m})=U_{\mathfrak{m}}$.
Note that for any maximal ideal $\mathfrak{m}$ of $R$, $U(\mathfrak{m})_{Q}=U_{Q}$ since all these modules have length $r$ over $\Lambda_{Q}$. Therefore, we can apply Theorem 5.10 and conclude that there exists a finitely generated $\Lambda$-module $P$ of $U$ such that $P(\mathfrak{m})\cong U(\mathfrak{m})=P_{\mathfrak{m}}$ for every $\mathfrak{m}\in\operatorname{mSpec}(R)$. By Corollary 5.5, $P$ is a projective module over $\Lambda$.
We introduce the notation necessary to more thoroughly detail the image of the genus map $\Phi$ and the rank map $\Psi$.
Notation 7.4.
@@
$\Phi\colon V^{*}(\Lambda)$$\prod_{\mathfrak{m}\in\operatorname{mSpec}(R)}V^{*}(\Lambda_{\mathfrak{m}})\cong\prod_{\mathfrak{m}\in\operatorname{mSpec}(R)}{\mathbb{N}}_{0}^{*}$}
@@
For $i=1,\dots,\ell$, let $U_{i,1},\dots,U_{i,t_{i}}$ be representatives of indecomposable projective right $\Lambda_{\mathfrak{m}_{i}}$-modules, and let $r_{i,1},\dots,r_{i,t_{i}}$ be the lengths of their localizations over $\Lambda_{Q}$, i.e., $(U_{i,j})_{Q}\cong S^{r_{i,j}}$.
Note that, for any $i=1,\dots,\ell$, if the right $\Lambda_{\mathfrak{m}_{i}}$-module $\Lambda_{\mathfrak{m}_{i}}$ is isomorphic to $U_{i,1}^{m_{i,1}}\oplus\dots\oplus U_{i,t_{i}}^{m_{i,t_{i}}}$, then we have the relation
@@
We proceed to construct examples of finitely generated locally semiperfect torsion-free $\mathbb{Z}$-algebras. In particular, they are algebras over an $h$-local domain with an infinite number of maximal ideals. These examples show that there are no other restrictions on the parameters $k,\ell,t_{i},r_{ij},m_{ij}$ than the $\ell$ equations above.
@@
    $\Lambda$ is a subring of ${\rm M}_{k}(\mathbb{Q})$, is a finitely generated $\mathbb{Z}_{(p)}$-algebra and $\Lambda_{Q}={\rm M}_{k}(\mathbb{Q})$.
@@
From $(iii)$ it follows that $\Lambda/J$ is a direct sum of simple modules. Hence $\Lambda/J$ is semisimple artinian and therefore $J=J(\Lambda)$. Also note that the canonical projection $e_{i,j,j}\Lambda\to e_{i,j,j}\Lambda/e_{i,j,j}J$ is a projective cover of $e_{i,j,j}\Lambda/e_{i,j,j}J$. So $\Lambda$ is semiperfect and $e_{i,j,j}\Lambda$ are indecomposable projectives. Also $e_{i,a,a}\Lambda\cong e_{j,b,b}\Lambda$ if and only if $i=j$. So we have that $\Lambda=\bigoplus_{i=1}^{t}\bigoplus_{a=1}^{m_{i}}e_{i,a,a}\Lambda\cong\bigoplus_{i=1}^{t}(e_{i,1,1}\Lambda)^{m_{i}}$. Also, it is easy to check that $e_{i,a,a}{\rm M}_{k}(\mathbb{Q})$ has length $r_{i}$ since $|A_{i,a}|=r_{i}$.
@@
Now the Package Deal Theorem (here it is enough to use the noetherian version from [levyodenthal2]) finds a finitely generated $\mathbb{Z}$-submodule $\Lambda\subseteq{\rm M}_{k}(\mathbb{Z})$ such that $\Lambda_{(p_{i})}=\Lambda(p_{i})$ for $i=1,\dots,\ell$ and $\Lambda_{(q)}={\rm M}_{k}(\mathbb{Z}_{(q)})$ if $q$ is a prime different from $p_{1},\dots,p_{\ell}$. It is easy to see that $\Lambda$ has to be a subring of ${\rm M}_{k}(\mathbb{Z})$ and that $\Lambda_{(p_{i})}\simeq\Lambda(p_{i})$ as rings.
@@
We keep the Notation 7.4. First, we observe that
$\Phi\colon V^{*}(\Lambda)$$\prod_{\mathfrak{m}\in\operatorname{mSpec}(R)}V^{*}(\Lambda_{\mathfrak{m}})=\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})\times\prod_{\mathfrak{m}\not\in\mathcal{S}}V^{*}(\Lambda_{\mathfrak{m}})$}
So, for any $\mathfrak{m}\not\in\mathcal{S}$, $V^{*}(\Lambda_{\mathfrak{m}})\cong{\mathbb{N}}_{0}^{*}$, and for any $P]\in V^{*}(\Lambda)$, this isomorphism sends $P_{\mathfrak{m}}]$ to its rank, which is $\Psi([P])$. As we are assuming that $\ell\geq 1$, the genus map is completely determined by its image in $\prod_{i=1}^{\ell}V^{*}(\Lambda_{\mathfrak{m}_{i}})$.
Then, there is a monoid morphism
@@
Where the isomorphism $\prod_{i=1}^{\ell}V(\Lambda_{\mathfrak{m}_{i}})\cong\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ assigns to $\Phi^{\prime}([P])$ the element
@@
Conversely, by Proposition 7.3, an element $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ corresponds to a finitely generated projective right $\Lambda$-module if and only if there exists $r$ such that the components of $\mathbf{X}$ satisfy the system of equations (10). This describes all the possible genus of finitely generated projective $\Lambda$-modules.
In summary, we have proved the following.
@@
Let $\mathcal{A}=\Phi^{\prime}(V(\Lambda))$ which is a submonoid of $\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$. Then $\mathbf{X}\in\prod_{i=1}^{\ell}\left({\mathbb{N}}_{0}^{t_{i}}\right)$ is an element of $\mathcal{A}$ if and only if there exists $r\in{\mathbb{N}}_{0}$ such that the components of $\mathbf{X}$ satisfy
@@
If $\ell=1$ and $U_{1},\dots,U_{t}$ are the indecomposable projective modules of $\Lambda_{\mathfrak{m}}$, up to isomorphism, and $\Psi(U_{i})=r_{i}$ for $i=1,\dots,t$. Then $\Psi(V(\Lambda))$ is the submonoid of ${\mathbb{N}}_{0}$ generated by $r_{1},\dots,r_{\ell}$. By Proposition 7.3, $\Phi^{\prime}(V(\Lambda))\cong{\mathbb{N}}_{0}^{t}$, because for any $i=1,\dots,\ell$, there exists an (indecomposable) finitely generated projective right $\Lambda$-module $P_{i}$ such that $(P_{i})_{\mathfrak{m}_{i}}\cong U_{i}$.
@@
$\sum_{j=1}^{t_{i}}r_{i,j}y_{i,j}=(r_{i,1},\dots,r_{i,t_{i}})\begin{pmatrix}y_{i,1}\\ \vdots\\ y_{i,t_{i}}\end{pmatrix}$$\infty$ (11)
@@
Let $\mathbf{Y}\in\prod_{i=1}^{\ell}\left(({\mathbb{N}}_{0}^{*})^{t_{i}}\right)$. For $i=1,\dots,\ell$, let $S_{i}(\mathbf{Y})=\{j\mid y_{i,j}=\infty\}$. Observe that $\mathbf{Y}$ is a solution of the equations (11) if and only if $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$.
Let $P$ be a countably generated right $\Lambda$-module such that it is not finitely generated. Then $\Psi(P)=\infty$, which implies that, for any $i=1,\dots,\ell$, there exists $j$ such that the indecomposable module $U_{i,j}$ appears infinitely many times in the decomposition of $P_{\mathfrak{m}}$ as a direct sum of indecomposable modules. Therefore, if $\mathbf{Y}=\Phi^{\prime}([P])$, then $S_{i}(\mathbf{Y})\neq\emptyset$ for any $i=1,\dots,\ell$. So $\mathbf{Y}$ is a solution to the system (11).
@@
For $i=1,\dots,\ell$, let $P^{\prime}(\mathfrak{m}_{i})=\bigoplus_{j\in S_{i}(\mathbf{Y})}U_{i,j}$ and let $r_{i}$ be the rank of $P^{\prime}(\mathfrak{m}_{i})$. Let $r$ be the least common multiple of $r_{1},\dots,r_{\ell},k$. For $i=1,\dots,\ell$, set $P(\mathfrak{m}_{i})=P^{\prime}(\mathfrak{m}_{i})^{\frac{m}{r_{i}}}$ and set $P(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{\frac{r}{k}}$ if $\mathfrak{m}$ is a maximal ideal of $R$ different from $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$. By Package deal Theorem 5.13, there exists a finitely generated projective $\Lambda$-module $P^{\prime}$ such that $P^{\prime}_{\mathfrak{m}}\cong P(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$.